### v1

@@ Abstract
The emergence of intelligent matter has sparked significant interest in next-generation technologies. We report on the discovery of a reconfigurable magnetic multilayer domain structure in the van der Waals magnet CrSBr, exhibiting a unique combination of magnetic and optical properties. Applying an external magnetic field along the easy axis drives the hysteretic antiferromagnetic–to–ferromagnetic transition that is not universally binary, but instead develops through a cascade of intermediate magnetic configurations whose multiplicity and stability scale systematically with thickness. This material can be considered as a prototypical intelligent matter, capable of encoding, processing, and storing information through its tunable magnetic structure. The directly linked optical properties of CrSBr, modulated by the magnetic structure, provide a readout mechanism for the stored information compatible with modern information distribution using light. With its adaptive properties, CrSBr is an attractive candidate for neuromorphic circuitries, enabling the design of brain-inspired computing architectures that can learn and evolve in response to changing environments.
@@ 1 Introduction
Reconfigurability, together with hysteresis, meaning that the magnetization in a magnet depends on its history, is attractive for information technologies enabling writing and storage of information. In this regard, antiferromagnets (AFM) offer ultrafast dynamics and multistability, while ferromagnets (FM) provide binary states but may suffer from magnetostatic (dipolar) crosstalk [8, 9, 10]. Encoding information in optically accessible AFM-derived states therefore opens pathways towards memory functionalities, spin-driven optoelectronics or even neuromorphic architectures. The latter requires that a system is capable to learn, process, store and distribute information by interacting with its environment, a combination of functionalities often referred to as ‘intelligent matter’ [11]. While writing and storage of digital information is at the heart of a magnetic systems with long-range order, learning capabilities, distribution of information and universal readout are often challenging.
@@
With a direct bandgap of around 1.3 eV and an unusually high refractive index near its excitonic resonances, even tens-of-nanometers thick flakes exhibit pronounced interference effects dictated by van der Waals stacking rather than external nanofabrication [25, 26, 27]. Here, we establish CrSBr as a naturally reconfigurable magneto-optical resonator enabling optical readout of a layer-by-layer magnetic domain structure that is controlled by external stimuli. The combined experimental and simulation results highlight the ‘intelligent matter’ capability of the vdW magnetic semiconductor CrSBr.
@@ Acknowledgements.
The authors gratefully acknowledge financial support by the German Science Foundation (DFG) via Grants No.443274199 and 556436549 (WU 637/7-2,8-1), and the priority program 2244 (2DMP) through start-up funding. This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101130224 ’JOSEPHINE’. T.D. acknowledges financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project No. 426726249 (DE 2749/2-1 and DE 2749/2-2). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) [42]. Z.S. was supported by project LUAUS25268 from the Ministry of Education, Youth and Sports (MEYS) and by the project Advanced Functional Nanorobots (reg. No. CZ.02.1.01/0.0/0.0/15__003/0000444 financed by the EFRR) and by the ERC-CZ program (project LL2101) from Ministry of Education, Youth and Sports (MEYS). Z.S. acknowledges the assistance provided by the Advanced Multiscale Materials for Key Enabling Technologies project, supported by the Ministry of Education, Youth, and Sports of the Czech Republic. Project No. CZ.02.01.01/00/22__008/0004558, Co-funded by the European Union.
@@
Figure S1: Comparison between three different approaches to the construction of the dielectric function used in the calculations, along with the subsequent result of the model. In each column: the top row shows the dispersion of the refractive index $n(E)$; the middle row shows the energy of the excitonic resonance depending on the number of layers, all other than energy parameters of the excitonic resonance used to the construction of the dielectric function are the same. The bottom row shows the normalized reflectance $R/R_{0}$ for exemplary configurations of AFM/FM stacks with different FM content (given in %). First approach (a) is differing only the monolayer exciton energy (AFM state) and FM state (the number of layers $\geqslant 2$). Even with this simple model we can observe characteristics of the intermediate AFM/FM stack — occurrence of more features in the spectrum and blueshifts. (b) The second approach with additional intermediate energy for thin layers (2-3L); The approach presented in (c) was used in the main paper.

### v2

@@ Abstract
The emergence of intelligent matter has sparked significant interest in next–generation technologies. We report on the discovery of a reconfigurable magnetic multilayer domain structure in the van der Waals magnet CrSBr, exhibiting a unique combination of magnetic and optical properties. Applying an external magnetic field along the easy axis drives the hysteretic antiferromagnetic–to–ferromagnetic transition that is not universally binary, but instead develops through a cascade of intermediate magnetic configurations whose multiplicity and stability scale systematically with thickness. This material can be considered as a prototypical intelligent matter, capable of encoding, processing, and storing information through its tunable magnetic structure. The directly linked optical properties of CrSBr, modulated by the magnetic structure, provide a readout mechanism for the stored information compatible with modern information distribution using light. With its adaptive properties, CrSBr is an attractive candidate for neuromorphic circuitries, enabling the design of brain-inspired computing architectures that can learn and evolve in response to changing environments.
@@ 1 Introduction
Reconfigurability, together with hysteresis, meaning that the magnetization in a magnet depends on its history, is attractive for information technologies enabling writing and storage of information. In this regard, antiferromagnets (AFM) offer ultrafast dynamics and multistability, while ferromagnets (FM) provide binary states but may suffer from magnetostatic (dipolar) crosstalk [8, 9, 10]. Encoding information in optically accessible AFM-derived states therefore opens pathways towards memory functionalities, spin-driven optoelectronics or even neuromorphic architectures. The latter requires that a system is capable to learn, process, store and distribute information by interacting with its environment, a combination of functionalities often referred to as “intelligent matter” [11]. While writing and storage of digital information is at the heart of a magnetic systems with long-range order, learning capabilities, distribution of information and universal readout are often challenging.
@@
With a direct bandgap of around 1.3 eV and an unusually high refractive index near its excitonic resonances, even tens-of-nanometers thick flakes exhibit pronounced interference effects dictated by van der Waals stacking rather than external nanofabrication [24, 25, 26]. Here, we establish CrSBr as a naturally reconfigurable magneto-optical resonator enabling optical readout of a layer-by-layer magnetic domain structure that is controlled by external stimuli. The combined experimental and simulation results highlight the “intelligent matter” capability of the vdW magnetic semiconductor CrSBr.
@@ Acknowledgements.
The authors gratefully acknowledge financial support by the German Science Foundation (DFG) via Grants No. 443274199 and 556436549 (WU 637/7-2,8-1), and the priority program 2244 (2DMP) through start-up funding. This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No. 101130224 “JOSEPHINE”. T.D. acknowledges financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project No. 426726249 (DE 2749/2-1 and DE 2749/2-2). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) [41]. Z.S. was supported by project LUAUS25268 from the Ministry of Education, Youth and Sports (MEYS) and by the project Advanced Functional Nanorobots (reg. No. CZ.02.1.01/0.0/0.0/15__003/0000444 financed by the EFRR) and by the ERC-CZ program (project LL2101) from Ministry of Education, Youth and Sports (MEYS). Z.S. acknowledges the assistance provided by the Advanced Multiscale Materials for Key Enabling Technologies project, supported by the Ministry of Education, Youth, and Sports of the Czech Republic. Project No. CZ.02.01.01/00/22__008/0004558, Co-funded by the European Union.
@@
Figure S1: Comparison between three different approaches to the construction of the dielectric function used in the calculations, along with the subsequent result of the model. In each column: the top row shows the dispersion of the refractive index $n(E)$; the middle row shows the energy of the excitonic resonance depending on the number of layers, all other than energy parameters of the excitonic resonance used to the construction of the dielectric function are the same. The bottom row shows the normalized reflectance $R/R_{0}$ for exemplary configurations of AFM/FM stacks with different FM content (given in %). First approach (a) is differing only the monolayer exciton energy (AFM state) and FM state (the number of layers $\geqslant 2$). Even with this simple model we can observe characteristics of the intermediate AFM/FM stack — occurrence of more features in the spectrum and blueshifts. (b) The second approach with additional intermediate energy for thin layers (2–3L); The approach presented in (c) was used in the main paper.